# Adiabatic hydrodynamization and quasinormal modes of nonthermal attractors

### v1

@@ Abstract
Nonthermal attractors characterize the emergent self-similar evolution of far-from-equilibrium quantum systems, from nuclear collisions to cold-atom experiments. Within the adiabatic hydrodynamization framework, the approach to the nonthermal attractor can be understood as the decay of excited states of an effective Hamiltonian. We use an exactly solvable case of the Boltzmann equation—the longitudinally expanding overoccupied gluon plasma dominated by small-angle elastic collisions—to map the eigenmodes in adiabatic hydrodynamization to the quasinormal mode spectrum of the nonthermal attractor. If generalized, this equivalence may be used as a discovery tool for new phenomena in out-of-equilibrium systems. A byproduct of our analysis is the analytic prescaling solutions for systems undergoing strong longitudinal expansion.
@@
$f(\tau,p_{z},p_{\tau})=\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\alpha}f_{s}\Big(\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\gamma}p_{z},\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\beta}p_{\perp}\Big)\,$, (1)
@@
On the other hand, the notion of QNMs has only recently been established in the context of non-expanding, isotropic NTFPs [21]. However, near equilibrium, the notion of QNMs is widespread in usage. In particular, in holographic systems black hole QNMs dictate how a small perturbation to the equilibrium state relaxes back to equilibrium [44, 45, 46] — their analogs in weakly coupled systems are determined by the linearized time evolution operator appearing in the Boltzmann equation. The QNMs of NTFPs provide a different perspective on the emergence and approach to universal out-of-equilibrium dynamics. In this case, they represent self-similar power law corrections to the NTFP solution Eq. (1). If the power laws have negative exponents, the QNMs decay in time, explaining the attractive nature of a NTFP.
@@
In this framework, the organizing principle is adiabaticity: only if the evolution is adiabatic, the low-lying eigenstates are guaranteed to be stable solutions to which the system will be attracted. While adiabaticity is not guaranteed a priori, previous works have shown [18, 19, 20] that it is often possible to obtain an adiabatic description of the dynamics by introducing suitable time-dependent coordinate redefinitions. In particular, Eq. (2) was solved in [18] by introducing arbitrary functions A,B,G{A},{B},{G} that rescale the distribution function and momenta in time: $f(\tau,p_{z},p_{\perp})={A}(\tau){\tilde{f}}(\tau,{\xi},{\chi})$, with ${\xi}\equiv{G}(\tau)p_{z}$ and ${\chi}\equiv{B}(\tau)p_{\perp}$. In terms of the rescaled variables, the kinetic equation (2) takes the form
@@
In this way, the non-linear partial differential equation (2) has been reduced to finding A,G{A},{G} that satisfy Eqs. (4) and (5). These in turn enter the eigenvalues and eigenstates above. Since $\tilde{I}$} depends on the initial conditions for all coefficients, a general closed form solution to these equations seems beyond reach. However, this can be achieved in certain cases, which we now examine.
@@
Eqs. (15), (18) and (19) define the prescaling solution of Eq. (2). In the following we will refer to $\bar{A}(\tau)^{-1}{f}(\tau,\bar{\xi},\bar{\chi})$ as the prescaling frame. In this frame the system becomes effectively static when it is in the prescaling regime. Although beyond our scope, the same solution could be applied to prescaling in the full EKT [49], though the scaling-breaking terms add complexity [48].
QNMs in Bjorken flow.– In [21] it was shown that, near a non-expanding, isotropic NTFP, linear order deviations from it can be written as a superposition of quasinormal modes (QNMs). These decay in a power-law manner to the NTFP with different exponents $\Omega$, which define the QNM frequencies. We generalize this analysis to the expanding system by considering a perturbation in the prescaling frame $\bar{A}(\tau)^{-1}f(\tau,p_{z},p_{\perp})=f_{s}(\bar{\xi},\bar{\chi})+\delta\bar{f}(\tau,\bar{\xi},\bar{\chi})$ with $\delta\bar{f}\ll 1$. At linear order in $\delta\bar{f}$, Eq. (2) yields an equation of the form
@@
where $\bar{I}_{\delta}=1/(2\pi)^{3}\int d^{2}\bar{\chi}\int d\bar{\xi}2f_{s}f_{\Omega}$. For simplicity, we will in the following assume all perturbations to have the same transverse momentum distribution as the background, which we alleviate in App. A. Eq. (23) can be solved analytically, resulting in the following QNM frequencies and modes.
@@
$\begin{split}&f_{\Omega_{1}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\\ &f_{\Omega_{n}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\Bigg[\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})\\ &\qquad\qquad+\frac{\bar{I}_{\text{He}_{2n}}}{\sqrt{\pi}\sqrt{\kappa}(\Omega_{n}-3)}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\Bigg],\ n\neq 1.\end{split}$ (25)
@@
For our solution of the scaling function Eq. (19), this indeed results in the QNM solution with $\Omega_{1}=3$, see Eq. (25).
The analogous analysis for $\tau_{\rm ref}\rightarrow\tau_{\rm ref}+\delta\tau_{\rm ref}$ leads to the QNM solution with $\Omega_{0}=0$. Note that in this case one needs to take into account the explicit dependence of $f_{s}$ on $\tau_{\rm ref}$. This enters through $\kappa\propto\tau_{\text{ref}}^{2}$, see Eq. (20).
@@
Similar nonlinear solutions have been considered in the context of black hole QNMs [50, 51]. There transient effects are studied, based on the fact that the QNMs are the eigenvalues of a non-normal operator. We anticipate that aspects of these findings could apply in the context of NTFP QNMs. To this end, a more careful investigation of the non-normality of the linearized operator in Eq. (23) is necessary.
@@
Now, by means of Eqs. (4) and (5), we are able to calculate A{A} as a series expansion in $\delta_{n}$: ${A}=\bar{A}+\sum_{i}^{\infty}(\delta_{n})^{i}~\bar{A}_{i}$. We will consider here the case $n=2$. Plugging the resulting solution into ${\tilde{f}}=\phi_{0}^{R}({\xi})+\delta_{2}\bar{A}(\tau)^{2}\phi^{R}_{2}({\xi})$ and truncating after $O(\delta_{2}^{2})$, we find
$\begin{split}{\tilde{f}}&=\bar{g}(\bar{\chi})\Bigg[e^{-\bar{\xi}^{2}/2}-(d_{1}+d_{2}\delta_{2})\delta_{2}\bar{A}(\tau)^{3}e^{-\bar{\xi}^{2}/2}\left(-1+\bar{\xi}^{2}\right)\\ &+\frac{1}{24}\delta_{2}\bar{A}(\tau)^{4}e^{-\bar{\xi}^{2}/2}\left(18-30\bar{\xi}^{2}+4\bar{\xi}^{4}\right)\\ &+\frac{1}{2}d_{1}^{2}\delta_{2}^{2}\bar{A}(\tau)^{6}e^{-\bar{\xi}^{2}/2}\left(4-7\bar{\xi}^{2}+\bar{\xi}^{4}\right)\\ &-\frac{1}{12}d_{1}\delta_{2}^{2}\bar{A}(\tau)^{7}e^{-\bar{\xi}^{2}/2}\left(-45+114\bar{\xi}^{2}-33\bar{\xi}^{4}+2\bar{\xi}^{6}\right)\\ &+\frac{1}{960}\delta_{2}^{2}\bar{A}(\tau)^{8}e^{-\bar{\xi}^{2}/2}\left(803-2093\bar{\xi}^{2}+630\bar{\xi}^{4}-40\bar{\xi}^{6}\right)\Bigg],\end{split}$ (32)
@@
where we made use of ${\xi}={G}(\tau)p_{z}$ and Eq. (4). $d_{1}$ and $d_{2}$ are integration constants that are picked up order by order in the solution for A{A}.
Comparing this to Eq. (30), we find that there is an exact match when considering terms to $O(\delta_{2}^{2})$ after identifying $c_{1}=-(d_{1}+d_{2}\delta_{2})\delta_{2}$ and $c_{2}=\delta_{2}/24$. This serves as a non-trivial cross-check of the earlier solution in Eq. (30), since this approach is agnostic to the expressions for the QNMs. More importantly, it is now clear how the QNMs enter the AH solution. The first excited state excites the second QNM $\Omega_{2}=4$ and the prescaling QNM $\Omega_{1}=3$. Notably, no contributions from other QNMs enter the solution. This structure persists when considering additional excited states on top of the ground state. The effect of the $n$th excited state can be captured by the prescaling and the (n1)(n-1)th QNMs.
Therefore, we achieved our goal of understanding the explicit time dependence of A{A} close to the NTFP. Indeed the approach is captured by a series of terms that decay according to different power-laws. This is the QNM picture we presented before. Importantly, the slowest power law is $\bar{A}^{3}$ instead of $\bar{A}^{2}$, which one might have naively expected from the AH picture. From the AH perspective, this is due to the fact that perturbations in the evolution of $A$ in terms of $\delta_{n}$, in addition to the perturbations themselves, generate a change to the coordinate frame on which they propagate, leading to a power law that is not directly determined by the eigenvalues $\mathcal{E}_{n}$. Most importantly, we have established the equivalence between the two pictures and how to construct the mapping.
@@
Outlook.– In this Letter, we have found explicit solutions for universal features of the dynamics of nonlinear systems, as well as organizing principles with which to generalize these findings. These give us a platform to develop concrete, rigorous way of understanding out-of-equilibrium phenomena. Specifically, by connecting the explicit QNM solutions around a NTFP to the exact (albeit implicit) nonlinear solutions of the dynamics of the same theory, we have opened the gates to a systematic exploration of the nonlinear response regime around NTFPs and the general relaxation dynamics towards them.
Our results explicitly show how to extend QNMs into the nonlinear regime. Furthermore, we have shown how these solutions can be matched to fully nonlinear solutions obtained in the AH framework. In tandem, these two developments show how one may seek and find exact adiabatic solutions that extend non-perturbatively away from the NTFP by using the information encoded in the QNMs. Here, we have focused on the BMSS fixed point, but we speculate that the steps we took to connect these two concepts may be applicable for a broad class of NTFPs. Establishing the breadth of this class — which our work exemplifies — is an exciting new open problem, with potential to give us deep insights about the general classification and structure of universal out-of-equilibrium phenomena.
@@
In this work, we have considered a simple idealized model, where a large anisotropy justifies the complete decoupling of the transverse sector. A natural next step would be to study how our results extend to models that go beyond our simplifying assumptions by including transverse dynamics or even fully dropping the assumption of longitudinal momentum domination. As well as for models that have a dual cascade [52, 53], making the connection stronger to experimental NTFPs that observe such dual cascades [28, 26, 24, 25, 27, 29, 30]. On an even broader scope, as the concept of QNMs originates from holography, recent insights about their role in that context [54, 55, 56, 57, 51, 50] might synergize with the results of this paper to gain an even deeper understanding of the process. Similarly, important insights might be gained from exploring a connection to singulants [58], which were introduced as an extension of the notion of QNMs beyond near-equilibrium systems, motivated by the large order behavior of the factorially divergent hydrodynamic gradient expansion.

### v2

@@ Abstract
Nonthermal attractors govern the emergent self-similar dynamics of far-from-equilibrium quantum systems, from ultrarelativistic nuclear collisions to cold-atom experiments. Within the framework of adiabatic hydrodynamization, the approach to a nonthermal attractor is described by the decay of excited states of an effective Hamiltonian. Using an exactly solvable kinetic theory—the longitudinally expanding, overoccupied gluon plasma dominated by small-angle elastic scattering—we establish a direct correspondence between the eigenmodes of adiabatic hydrodynamization and the quasinormal mode spectrum of the nonthermal attractor. This equivalence suggests a general framework for identifying universal dynamical structures in nonequilibrium systems. As a byproduct, we derive analytic prescaling solutions for strongly longitudinally expanding systems.
@@
$f(\tau,p_{z},p_{\perp})=\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\alpha}f_{s}\Big(\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\gamma}p_{z},\big(\tfrac{\tau}{\tau_{\text{ref}}}\big)^{\beta}p_{\perp}\Big)\,$, (1)
@@
On the other hand, the notion of QNMs has only recently been established in the context of non-expanding, isotropic NTFPs [21]. However, near equilibrium, this notion is widespread in usage. In particular, in holographic systems black hole QNMs dictate how a small perturbation to the equilibrium state relaxes back to equilibrium [44, 45, 46] — their analogs in weakly coupled systems are determined by the linearized time evolution operator appearing in the Boltzmann equation. The strength of a QNM analysis is that it is based on a linearization around any given state. This means the procedure is universal and can in principle always be performed. The QNMs of NTFPs provide a different perspective on the emergence of and approach to universal out-of-equilibrium dynamics. In this case, they represent self-similar power law corrections to the NTFP solution Eq. (1). If the power laws have negative exponents, the QNMs decay in time, explaining the attractive nature of a NTFP.
@@
In this framework, the organizing principle is adiabaticity: only if the evolution is adiabatic, are the low-lying eigenstates guaranteed to be stable solutions to which the system will be attracted. While adiabaticity is not guaranteed a priori, previous works have shown [18, 19, 20] that it is often possible to obtain an adiabatic description of the dynamics by introducing suitable time-dependent coordinate redefinitions. In particular, Eq. (2) was solved in [18] by introducing arbitrary functions A,B,G{A},{B},{G} that rescale the distribution function and momenta in time: $f(\tau,p_{z},p_{\perp})={A}(\tau){\tilde{f}}(\tau,{\xi},{\chi})$, with ${\xi}\equiv{G}(\tau)p_{z}$ and ${\chi}\equiv{B}(\tau)p_{\perp}$. In terms of the rescaled variables, the kinetic equation (2) takes the form
@@
In this way, the non-linear partial differential equation (2) has been reduced to finding A,G{A},{G} that satisfy Eqs. (4) and (5). We call this pair A,G{A},{G} the adiabatic frame. These in turn enter the eigenvalues and eigenstates above. Since $\tilde{I}$} depends on the initial conditions for all coefficients, a general closed form solution to these equations seems beyond reach. However, this can be achieved in certain cases, which we now examine.
@@
Eqs. (15), (18) and (19) define the prescaling solution of Eq. (2). In the following we will refer to $\bar{A}(\tau)^{-1}{f}(\tau,\bar{\xi},\bar{\chi})$ as the prescaling frame. In this frame the system becomes effectively static when it is in the prescaling regime. Although beyond our scope, the same ansatz could be applied to prescaling in the full EKT [49], though the scaling-breaking terms add complexity [48].
QNMs in an expanding NTFP.– In [21] it was shown that, near a non-expanding, isotropic NTFP, linear order deviations from it can be written as a superposition of quasinormal modes (QNMs). These decay in a power-law manner to the NTFP with different exponents $\Omega$, which define the QNM frequencies. We generalize this analysis to the expanding system by considering a perturbation in the prescaling frame $\bar{A}(\tau)^{-1}f(\tau,p_{z},p_{\perp})=f_{s}(\bar{\xi},\bar{\chi})+\delta\bar{f}(\tau,\bar{\xi},\bar{\chi})$ with $\delta\bar{f}\ll 1$. At linear order in $\delta\bar{f}$, Eq. (2) yields an equation of the form
@@
where $\bar{I}_{\delta}=1/(2\pi)^{3}\int d^{2}\bar{\chi}\int d\bar{\xi}2f_{s}f_{\Omega}$. For simplicity, we will in the following assume all perturbations to have the same transverse momentum distribution as the background, and discuss the general case in App. A. Eq. (23) can be solved analytically, resulting in the following QNM frequencies and modes.
@@
$\begin{split}&f_{\Omega_{1}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\\ &f_{\Omega_{n}}(\bar{\xi},\bar{\chi})\propto\bar{g}(\bar{\chi})e^{-{\bar{\xi}^{2}}/({2\kappa})}\Bigg[\text{He}_{2n}(\bar{\xi}/\sqrt{\kappa})\\ &\qquad\qquad-\frac{\bar{I}_{\text{He}_{2n}}}{\sqrt{\pi}\sqrt{\kappa}(\Omega_{n}-3)}\text{He}_{2}(\bar{\xi}/\sqrt{\kappa})\Bigg],\ n\neq 1.\end{split}$ (25)
@@
By comparing the spectra in the AH picture (Eq. (7)) and the QNM picture (Eq. (25)), we make the following observation: All right AH eigenstates are given by a linear combination of the corresponding QNM and the $\Omega=3$ mode. The latter is related to prescaling as shown in the next section. This observation, which we investigate further at the non-linear level in subsequent sections, gives a first glimpse at how the two frameworks are related.
@@
For our solution of the scaling function Eq. (19), this indeed results in the QNM solution with $\Omega_{1}=3$, see Eq. (25), in the prescaling frame $\bar{A}(\tau)^{-1}f$.
The analogous analysis for $\tau_{\rm ref}\rightarrow\tau_{\rm ref}+\delta\tau_{\rm ref}$ leads to the QNM solution with $\Omega_{0}=0$. In this case one needs to take into account the explicit dependence of $f_{s}$ on $\tau_{\rm ref}$. This enters through $\kappa\propto\tau_{\text{ref}}^{2}$, see Eq. (20).
@@
Similar nonlinear solutions have been considered in the context of black hole QNMs [50, 51]. There transient effects are studied, based on the fact that the QNMs are the eigenvalues of a non-normal operator. We anticipate that aspects of these findings could apply in the context of NTFP QNMs.
@@
Now, by means of Eqs. (4) and (5), we are able to calculate A{A} as a series expansion in $\delta_{n}$: ${A}=\bar{A}+\sum_{i}^{\infty}(\delta_{n})^{i}~\bar{A}_{i}$. We will consider here the case $n=2$. Plugging the resulting solution into ${\tilde{f}}=\phi_{0}^{R}({\xi})+\delta_{2}\bar{A}(\tau)^{4}\phi^{R}_{2}({\xi})$ and truncating after $O(\delta_{2}^{2})$, we find
$\begin{split}\bar{f}&=\bar{g}(\bar{\chi})\Bigg[e^{-\bar{\xi}^{2}/2}-(d_{1}+d_{2}\delta_{2})\delta_{2}\bar{A}(\tau)^{3}e^{-\bar{\xi}^{2}/2}\left(-1+\bar{\xi}^{2}\right)\\ &+\frac{1}{96}\delta_{2}\bar{A}(\tau)^{4}e^{-\bar{\xi}^{2}/2}\left(18-30\bar{\xi}^{2}+4\bar{\xi}^{4}\right)\\ &+\frac{1}{2}d_{1}^{2}\delta_{2}^{2}\bar{A}(\tau)^{6}e^{-\bar{\xi}^{2}/2}\left(4-7\bar{\xi}^{2}+\bar{\xi}^{4}\right)\\ &-\frac{1}{48}d_{1}\delta_{2}^{2}\bar{A}(\tau)^{7}e^{-\bar{\xi}^{2}/2}\left(-45+114\bar{\xi}^{2}-33\bar{\xi}^{4}+2\bar{\xi}^{6}\right)\\ &+\frac{1}{15360}\delta_{2}^{2}\bar{A}(\tau)^{8}e^{-\bar{\xi}^{2}/2}\left(803-2093\bar{\xi}^{2}+630\bar{\xi}^{4}-40\bar{\xi}^{6}\right)\Bigg],\end{split}$ (32)
@@
where we made use of ${A}(\tau){\tilde{f}}=f=\bar{A}(\tau)\bar{f}$ and ${G}(\tau)^{-1}{\xi}=p_{z}=\bar{G}(\tau)^{-1}\bar{\xi}$ together with Eq. (4). $d_{1}$ and $d_{2}$ are integration constants that are picked up order by order in the solution for A{A}.
Comparing this to Eq. (30), we find that there is an exact match when considering terms to $O(\delta_{2}^{2})$ after identifying $c_{1}=-(d_{1}+d_{2}\delta_{2})\delta_{2}$ and $c_{2}=\delta_{2}/96$. This serves as a non-trivial cross-check of the earlier solution in Eq. (30), since this approach is agnostic to the expressions for the QNMs. Therefore, we achieved our goal of understanding the explicit time dependence of A{A} close to the NTFP. It is indeed given by a series of power-laws following the QNM picture we presented before. Note that, contrary to what one may have expected from the AH picture, the slowest power law is $\bar{A}^{3}$ instead of $\bar{A}^{2}$.
More importantly, we have now proven that the correspondence between QNMs and AH modes we saw at the linear level (see text below Eq. (25)) also holds in the perturbative nonlinear regime. The second excited state in AH excites the second QNM $\Omega_{2}=4$ and the prescaling QNM $\Omega_{1}=3$. Notably, no contributions from other QNMs enter the solution. This structure persists when considering additional excited states on top of the ground state, where more generally the $n$th excited state can be described by the $n$th and the prescaling QNMs. This is so because the difference between the adiabatic and prescaling frames is that the scaling variables A{A} and G{G} in the former frame are sourced by the excited states, and thus take different values than in the prescaling frame. At the linear level, an equivalent change in the value of $\bar{A}$ and $\bar{G}$ could be achieved by adjusting the integration constants $\tau_{\rm ref}$ and $\tau_{*}$. However, in the QNM picture such a change is viewed not as changing $\bar{A}$ and $\bar{G}$, but rather as modifying the distribution function itself. This is merely a reflection of the fact that the definition of the scaling function $f_{s}$ can always be changed via constant rescalings of $\bar{A}$ and $\bar{G}$.
@@
Outlook.– In this Letter, we have found explicit solutions for the dynamics of a specific nonlinear system around its NTFP solution, both at the linear and nonlinear level. These give us a platform to develop concrete, rigorous ways of understanding out-of-equilibrium phenomena. Specifically, by connecting the explicit QNM solutions around a NTFP which we extended to the perturbative nonlinear regime to the exact, but implicit, nonlinear solutions of the dynamics of the same theory obtained via AH, we have opened the gates to a systematic exploration of the nonlinear response regime around NTFPs and the general relaxation dynamics towards them.
For instance, our work provides the foundation for future work in constructing a general algorithm one may employ to find exact adiabatic solutions that extend non-perturbatively away from the NTFP by using the information encoded in the QNMs. This algorithm would exploit the fact that a QNM analysis can always be applied close to a state of universal dynamics and leads to a classification of excitations near this state, which can serve as a basis for constructing the AH spectrum.
There are many models that would be interesting to explore using this approach. In this Letter, we have considered an idealized description of a boost-invariant longitudinally expanding gluonic plasma, where a large anisotropy justifies the complete decoupling of the transverse sector. This approximation allowed us to derive closed-form expressions for prescaling and it would be interesting to compare these to simulations of QCD EKT and see if our predictions capture the prescaling dynamics found in [10]. Another natural next step would be to study how our results extend to models that go beyond our simplifying assumptions by including transverse dynamics or even fully dropping the assumption of longitudinal momentum domination.
@@
On a broader scope, as the concept of QNMs originates from holography, recent insights about their role in that context [52, 53, 54, 55, 51, 50] might synergize with the results of this paper to gain an even deeper understanding of the process. Similarly, important insights might be gained from exploring a connection to singulants [56], which were introduced as an extension of the notion of QNMs beyond near-equilibrium systems, motivated by the large order behavior of the factorially divergent hydrodynamic gradient expansion.